OpenStudy (anonymous):

A sperical balloon is being blown up in such a way that it's volume increases at the constant rate of 2 cubic metres per minute. Find the rate at which the radius is increasing at the instant when it is 3 metres. (HINT- the formula for the volume of a sphere of radius r is V= 3/4 πr^3) Thanks in advance!

OpenStudy (campbell_st):

You seem to have the volum of a sphere written incorrectly its V = 4/3 pi r^3 isn't this is case of related rates. \[\frac{dV}{dt} = \frac{dV}{dr} \times \frac{dr}{dt} \] so \[\frac{dV}{dr} = 4 \pi r^2\] so \[2 = 4 \pi r^2 \times \frac{dr}{dt}\] or \[\frac{dr}{dt} = \frac{2}{2 \pi r^2} \] now just substitute r = 3 to find dr/dt

OpenStudy (campbell_st):

oops it should read \[\frac{dr}{dt} = \frac{2}{4 \pi r^2} \] substitute r = 3 into the right hand side of the equation to find dr/dt